OpenStudy (anonymous):

how do i find the median and mean from a table with the first column being height - 120 - 130, and then next 130-140 and so on and then the second column being frequency... just wandering if you could help cheers

OpenStudy (anonymous):

with a lack of information about the data points in each group, you will have to make an assumption for the middle of each group ( 125, 135, etc.) Median would be the middle value, so you would have to figure how many total entries there were,and then figure what group the middle value falls in. The mean would be similar, add them all together (assumed magnitude which is followed by a 5, like 125, 135) multiplied by the actual height of the bar. The mean is then the total of all the height of the bars multiplied by their middle value, all divided by the number of total entries.

OpenStudy (anonymous):

so is the mean: the mid point times frequency, divided by the total freuqncey and the median is...? is it the highest value in the frequency table?

OpenStudy (anonymous):

so you have a table with two columns, a bar chart type table? The median in that case would be the value that is half the number of total samples. For example if for 120-130 the frequency was 5, next was 8, next was 4, the median would be where the (5+8+4)/2=17/2=(8th value), so the median would be 135 in this case as it was in the second group

OpenStudy (anonymous):

ok thanks i think i understand